OpenStudy (anonymous):

What is the result of adding the system of equations? 2x + y = 4 3x - y = 6

jimthompson5910 (jim_thompson5910):

what is 2x+3x

OpenStudy (anonymous):

5x

OpenStudy (uri):

2x + y = 4 3x - y = 6 ======= 5x=10 x=5 +y and -y will be cancelled.

jimthompson5910 (jim_thompson5910):

what is y + (-y)

OpenStudy (anonymous):

-y?

jimthompson5910 (jim_thompson5910):

think of y + (-y) as y - y

jimthompson5910 (jim_thompson5910):

it's like saying 1 + (-1) is the same as 1 - 1

OpenStudy (anonymous):

oh so 0

jimthompson5910 (jim_thompson5910):

so y + (-y) = 0y which is basically 0

jimthompson5910 (jim_thompson5910):

finally, 4 + 6 = ??

OpenStudy (anonymous):

10

jimthompson5910 (jim_thompson5910):

that explains how adding the equations (all the like terms matching up and adding) gives you this 2x + y = 4 + 3x - y = 6 ---------------- 5x + 0y = 10 which turns into 5x + 0 = 10 5x = 10

jimthompson5910 (jim_thompson5910):

you then solve for x afterwards, you use that value of x you just found to find y

OpenStudy (anonymous):

Ok

jimthompson5910 (jim_thompson5910):

but you can stop at 5x = 10 because it just wants to know what you get when you add the equations

OpenStudy (anonymous):

But if i didnt have to stop would the awnser be 2x?

jimthompson5910 (jim_thompson5910):

well if you want the answer for x and y, you would keep going

jimthompson5910 (jim_thompson5910):

if 5x = 10, then what is x

OpenStudy (anonymous):

it would be 10/2 right?

jimthompson5910 (jim_thompson5910):

5x = 10 x = ??? and no it's not 10/2